# Spatial Spiking Neural Networks Enable Efficient and Robust Temporal Computation

### v1

@@ Abstract
The efficiency of modern machine intelligence is measured by a high model accuracy and a low computational effort. Specifically in spiking neural networks (SNNs), synaptic delays are essential for capturing temporal patterns, yet current approaches treat delays as fully trainable, unconstrained parameters which incur large memory footprints, increased computational cost, and a disconnect from biological reality. In contrast, delays in the biological brain arise from the physical distance between neurons embedded in space. Inspired by this principle, we introduce Spatial Spiking Neural Networks (SpSNNs), a general framework in which neurons learn positions in a finite-dimensional Euclidean space and communication delays emerge from inter-neuron distances. This replaces per-synapse delay learning with position learning, drastically reducing the number of trainable parameters while preserving temporal expressiveness. Across two benchmark tasks (Yin-Yang and Spiking Heidelberg Digits), SpSNNs outperform conventional SNNs with unconstrained delays despite using far fewer parameters. Contrary to intuition, we find that accuracy consistently peaks in 2D and 3D networks rather than in infinite-dimensional ones, revealing a beneficial regularization effect. On top of that, dynamically sparsified SpSNNs retain full accuracy even at 90% sparsity, achieving equal performance with up to 18 fewer parameters compared to standard, delay-trained SNNs. Given these outcomes and the fact that the spatial structure of SpSNNs can directly translate into geometrical placements, SpSNNs in principle lend themselves also to efficient hardware implementations. In terms of methodology, SpSNNs rely on calculating exact gradients for trainable delays by elegant automatic differentiation with novel, custom-derived gradients, thus allowing arbitrary neural models and network architectures. Thereby, our proposed SpSNNs provide a platform for exploring the spatial dimension of temporal expressivity in bio-inspired AI models and provide a hardware-friendly substrate for scalable, energy-efficient neuromorphic intelligence.
@@ Introduction
By grounding temporal computation in learnable spatial structures, SpSNNs offer a biologically inspired and theoretically principled foundation for efficient spiking-based machine intelligence. Beyond their conceptual contributions, in this work we show that SpSNNs also confer practical advantages for deployment. Representing delay structure through low-dimensional position embeddings rather than large delay matrices reduces memory footprint and computational overhead, enabling more efficient execution on conventional hardware platforms such as CPUs, GPUs and AI accelerators landsmeer2024tricking. Moreover, because learned neuron coordinates correspond directly to geometric placement, SpSNNs naturally map onto two- or three-dimensional neuromorphic chips, simplifying placement, routing and delay management in hardware implementations. These properties position SpSNNs as a substrate for scalable, energy-efficient spiking computation across both standard AI infrastructure and emerging neuromorphic systems.
@@ Results
For the YY task, the exact timing of the only spike that neurons emit carry information, which makes them quite sensitive compared to other coding schemes bonilla2022ttfs. Figure 2D shows that SpSNNs can achieve around 93% test accuracy, which is on par with the results achieved in related work goltz2025delgrad; goltz2021fast; Wunderlich2021eventprop. This is despite the fact that SpSNNs are in a disadvantage in terms of temporal resolution due to their time-discretized simulations. This shows that SpSNNs managed to learn the sensitive temporal requirements of the YY task, while maintaining their generality.
In contrast, the SHD task uses a rate-based encoding, which reduces, but does not eliminate, the importance of exact spike timing guo2021coding. We can see in Figure 3D that SpSNNs achieve around 84% test accuracy with 300 recurrent neurons, a performance comparable to prior work hammouamri2023learningdelays; cramer2020heidelberg; Mészáros2025efficient; sun2025attentional. Notably, these previous approaches rely on additional mechanisms such as attention-based network architectures sun2025attentional, event-based simulation Mészáros2025efficient, or special measures during training hammouamri2023learningdelays, whereas SpSNNs achieve similar results within a more unified and flexible framework.
@@ Acknowledgments
This paper is partially supported by the European-Union Horizon Europe R&I program through projects SEPTON (no. 101094901) and SECURED (no. 101095717) and through the NWO - Gravitation Programme DBI2 (no. 024.005.022). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10677 and EINF-15815. The RTX PRO 6000 and Quadro Pro 6000 GPUs used for this research were donated by the NVIDIA Corporation.

### v2

@@ Abstract
The efficiency of modern machine intelligence is measured by a high model accuracy and a low computational effort. Specifically in spiking neural networks (SNNs), synaptic delays are essential for capturing temporal patterns, yet current approaches treat delays as fully trainable, unconstrained parameters which incur large memory footprints, increased computational cost, and a disconnect from biological reality. In contrast, delays in the biological brain arise from the physical distance between neurons embedded in space. Inspired by this principle, we introduce Spatial Spiking Neural Networks (SpSNNs), a general framework in which neurons learn positions in a finite-dimensional Euclidean space and communication delays emerge from inter-neuron distances. This replaces per-synapse delay learning with position learning, drastically reducing the number of trainable parameters while preserving temporal expressiveness. Across two benchmark tasks (Yin-Yang and Spiking Heidelberg Digits), SpSNNs outperform conventional SNNs with unconstrained delays despite using far fewer parameters. Contrary to intuition, we find that accuracy consistently peaks in 2D and 3D networks rather than in infinite-dimensional ones, revealing a beneficial regularization effect. On top of that, dynamically sparsified SpSNNs retain their accuracy even at 90% sparsity, achieving equal performance with up to 18 fewer parameters compared to standard, delay-trained SNNs. Given these outcomes and the fact that the spatial structure of SpSNNs can directly translate into geometrical placements, SpSNNs in principle lend themselves also to efficient hardware implementations. In terms of methodology, SpSNNs rely on calculating exact gradients for trainable delays by elegant automatic differentiation with novel, custom-derived gradients, thus allowing arbitrary neural models and network architectures. Thereby, our proposed SpSNNs provide a platform for exploring the spatial dimension of temporal expressivity in bio-inspired AI models and provide a hardware-friendly substrate for scalable, energy-efficient neuromorphic machine intelligence.
@@ Introduction
By grounding temporal computation in learnable spatial structures, SpSNNs offer a biologically inspired and theoretically principled foundation for efficient spiking-based machine intelligence. Beyond their conceptual contributions, in this work we show that SpSNNs also confer practical advantages for deployment. Representing delay structure through low-dimensional position embeddings rather than large delay matrices reduces memory footprint and computational overhead, enabling more efficient execution on conventional hardware platforms such as CPUs, GPUs and AI accelerators landsmeer2024tricking. Moreover, because learned neuron coordinates correspond directly to geometric placement, SpSNNs naturally map onto two- or three-dimensional neuromorphic chips, simplifying placement, routing and delay management in hardware implementations. These properties position SpSNNs as a promising substrate for scalable, energy-efficient spiking computation across both standard AI infrastructure and emerging neuromorphic systems.
@@ Results
For the YY task, the exact timing of the only spike that neurons emit carry information, which makes them quite sensitive compared to other coding schemes bonilla2022ttfs. Figure 2D shows that SpSNNs can achieve up to 98.10.1% test accuracy, which is on par with the results achieved in related work goltz2025delgrad; goltz2021fast; Wunderlich2021eventprop. This is despite the fact that SpSNNs are in a disadvantage in terms of temporal resolution due to their time-discretized simulations. This shows that SpSNNs managed to learn the sensitive temporal requirements of the YY task, while maintaining their generality.
In contrast, the SHD task uses a rate-based encoding, which reduces, but does not eliminate, the importance of exact spike timing guo2021coding. We can see in Figure 3D that SpSNNs achieve up to 83.60.9% test accuracy with 300 recurrent neurons, a performance comparable to prior work hammouamri2023learningdelays; cramer2020heidelberg; Mészáros2025efficient; sun2025attentional. Notably, these previous approaches rely on additional mechanisms such as attention-based network architectures sun2025attentional, event-based simulation Mészáros2025efficient, or special measures during training hammouamri2023learningdelays, whereas SpSNNs achieve similar results within a more unified and flexible framework.
@@ Acknowledgments
This paper is partially supported by the European-Union Horizon Europe R&I program through projects SEPTON (no. 101094901) and SECURED (no. 101095717) and through the NWO - Gravitation Programme DBI2 (no. 024.005.022). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-10677, EINF-15815, and EINF-16791. The RTX PRO 6000 and Quadro Pro 6000 GPUs used for this research were donated by the NVIDIA Corporation.